OpenStudy (anonymous):

Find the following quotient. (Divide.) -60/-15

OpenStudy (anonymous):

divede by 3

OpenStudy (anonymous):

do the cancilations

OpenStudy (anonymous):

0k

OpenStudy (mathstudent55):

When you multiply or divide two numbers with the same sign (both positive or both negative) the answer is positive. Otherwise, the answer is negative. Divide the numbers ignoring the signs, then use the above rule to determine which sign to give the answer. For example: 32/8 = 4 -32/-8 = 4 32/-8 = -4 -32/8 = -4

OpenStudy (anonymous):

-20

OpenStudy (anonymous):

-15/3=-5

OpenStudy (mathstudent55):

What is 60/15?

OpenStudy (anonymous):

4

OpenStudy (mathstudent55):

-15/3 = -5 is correct.

OpenStudy (mathstudent55):

60/15 = 4 is correct, so -60/-15 = 4 is also correct.

OpenStudy (anonymous):

so if it two negatives it equals a positive?

OpenStudy (mathstudent55):

Since -60 and -15 are both negative, the result of the division is positive. Since 60/15 = 4, -60/-15 = 4.

OpenStudy (anonymous):

therefore the answer is 4

OpenStudy (mathstudent55):

Correct. This is the rule for multiplication and division only. When you multiply or divide two numbers with the same sign, the answer is positive. When you multiply or divide two numbers with different signs, the answer is negative.